OpenStudy (anonymous):

if sin^-1(x/5) + cosec^-1(5/4) = pie/2 x =?

OpenStudy (anonymous):

@satellite73

OpenStudy (shubhamsrg):

cosec^-1 x = sin^-1 (1/x) can you try now ?

OpenStudy (shubhamsrg):

also you may use sin^-1 x + cos^-1 x =pi/2

OpenStudy (anonymous):

No idea

OpenStudy (shubhamsrg):

sin^-1(x/5) + cosec^-1(5/4) = pi/2 =>sin^-1(x/5) + sin^-1(4/5) = pi/2 =>sin^-1(x/5) + sin^-1(4/5) = sin^-1(4/5) + cos^-1 (4/5) =>sin^-1(x/5) = cos^-1 (4/5) now try to compare with this triangle  you should be able to do now.. 2 notes : 1) you could have even solve theoritically without considering the triangle by letting sin^(x/5) = y and' x/5 = sin y and blah blah blah... 2)its "pi" and not "pie" <-both are different :P

OpenStudy (anonymous):

thxx

OpenStudy (experimentx):